# Initial-State Typicality in Quantum Relaxation

### v1

@@ Abstract
where we use $\|R_{2}\|_{1}=\|R_{2}^{\dagger}\|_{1}=1$. From Eq. (4), it is clear that the relaxation timescale is determined by the slowest mode, i.e., $\tau_{\rm rel}\sim 1/|{\rm Re}(\lambda_{2})|$. Eq. (4) also implies that, the relaxation starting from an initial state with a smaller $|a_{2}|$ will be faster than that from one with a larger $|a_{2}|$  [^1] This can be more clearly seen by defining an $\epsilon$-relaxation time $\tau^{\epsilon}_{\rm rel}$ as in [32]. It is defined by the relation $|a_{2}|e^{-|{\rm Re}(\lambda_{2})|\tau^{\epsilon}_{\rm rel}}=\epsilon$, which gives $\tau^{\epsilon}_{\rm rel}=\frac{1}{|{\rm Re}(\lambda_{2})|}\ln(|a_{2}|/\epsilon)$.. In particular, if $a_{2}=0$, the relaxation timescale becomes $1/|{\rm Re}(\lambda_{3})|$, termed the strong Mpemba effect (SME) [28]. While the slowest mode, which has been the sole focus of most previous studies, governs the long-time dynamics, at intermediate times other modes $(k>2)$ can also contribute, motivating a more complete view.
@@
However, fast relaxation is not a necessary condition for typicality. Since the closing of Liouvillian gap is sufficient for a slow relaxation, the slow relaxation alone cannot imply $O_{k}$ grows fast.
@@
For most Hamiltonians without long-range interactions, $\Delta E\sim\mathcal{O}(n)\equiv cn$, where $c$ is a size-independent constant and $n$ is the system size. This requirement reflects the extensivity condition, which can also be fulfilled by long-range Hamiltonians when the Kac prescription is applied [92]. Then, take the many-qubit system as an example, if the temperature satisfies the condition $e^{\beta c}\leq 2\Rightarrow\beta<\beta_{c}:=\ln 2/c$, i.e., higher than a threshold irrelevant to the system size, $O_{k}$ has a smaller scaling than $\mathcal{O}(\sqrt{d})$ and all $a_{k}$ will concentrate around their mean for an initial Haar ensemble. For the HS ensemble, the threshold turns to $\beta_{c}^{\prime}=2\ln 2/c$.
@@
Two important implications of the typicality—First, the Liouvillian gap $1/|{\rm Re}(\lambda_{2})|$ may not correctly quantify the relaxation time for a class of systems satisfying two conditions: (i) $a_{2}$ concentrates (ii) $|\langle a_{2}\rangle|=|{\rm tr}(L_{2}^{\dagger})|/d$, vanishes. The second condition is met, for example, if ${\rm tr}(L_{2})=0$ [e.g., for a complex $\lambda_{2}$ of a Davies map [30]]. When both conditions hold, $a_{2}$ concentrates sharply around 0. Our Example 1 also falls into this class. We term this the “typical SME” (TSME), as it implies the SME ($a_{2}\approx 0$) is typical for generic initial states. In this regime, the relevant relaxation timescale may be set by $1/|{\rm Re}(\lambda_{k})|$ with $k\geq 3$, instead of $1/|{\rm Re}(\lambda_{2})|$. We discuss the conditions for this phenomenon in detail in [82].

### v2

@@ Abstract
where we use $\|R_{2}\|_{1}=\|R_{2}^{\dagger}\|_{1}=1$. From Eq. (4), it is clear that the relaxation timescale is determined by the slowest mode, i.e., $\tau_{\rm rel}\sim 1/|{\rm Re}(\lambda_{2})|$. Eq. (4) also implies that, the relaxation starting from an initial state with a smaller $|a_{2}|$ will be faster than that from one with a larger $|a_{2}|$  [^1] This can be more clearly seen by defining an $\epsilon$-relaxation time $\tau^{\epsilon}_{\rm rel}$ as in [32]. It is defined by the relation $|a_{2}|e^{-|{\rm Re}(\lambda_{2})|\tau^{\epsilon}_{\rm rel}}=\epsilon$, which gives $\tau^{\epsilon}_{\rm rel}=\frac{1}{|{\rm Re}(\lambda_{2})|}\log(|a_{2}|/\epsilon)$.. In particular, if $a_{2}=0$, the relaxation timescale becomes $1/|{\rm Re}(\lambda_{3})|$, termed the strong Mpemba effect (SME) [28]. While the slowest mode, which has been the sole focus of most previous studies, governs the long-time dynamics, at intermediate times other modes $(k>2)$ can also contribute, motivating a more complete view.
@@
However, fast relaxation is not a necessary condition for typicality. Since the closing of Liouvillian gap is sufficient for a slow relaxation, the slow relaxation alone cannot imply $O_{k}$ grows fast. Notably, a recent work provided an example for the typicality in slow relaxation [88] with a different normalization $\|R_{k}\|_{2}=1$. In their case, $\|L_{2}\|_{2}\approx\|\mathbb{I}_{d}\|=\sqrt{d}$ and thus ${\rm Var}(a_{2})\sim d^{-1}$.
@@
For most Hamiltonians without long-range interactions, $\Delta E\sim\mathcal{O}(n)\equiv cn$, where $c$ is a size-independent constant and $n$ is the system size. This requirement reflects the extensivity condition, which can also be fulfilled by long-range Hamiltonians when the Kac prescription is applied [93]. Then, take the many-qubit system as an example, if the temperature satisfies the condition $e^{\beta c}\leq 2\Rightarrow\beta<\beta_{c}:=\log 2/c$, i.e., higher than a threshold irrelevant to the system size, $O_{k}$ has a smaller scaling than $\mathcal{O}(\sqrt{d})$ and all $a_{k}$ will concentrate around their mean for an initial Haar ensemble. For the HS ensemble, the threshold turns to $\beta_{c}^{\prime}=2\log 2/c$.
@@
Two important implications of the typicality—First, the Liouvillian gap $1/|{\rm Re}(\lambda_{2})|$ may not correctly quantify the relaxation time for a class of systems satisfying two conditions: (i) $a_{2}$ concentrates (ii) $|\langle a_{2}\rangle|=|{\rm tr}(L_{2}^{\dagger})|/d$, vanishes. The second condition is met, for example, if ${\rm tr}(L_{2})=0$ [e.g., for a complex $\lambda_{2}$ of a Davies map [30]]. When both conditions hold, $a_{2}$ concentrates sharply around 0. Our Example 1 also falls into this class. We term this the “typical SME” (TSME), as it implies the SME ($a_{2}\approx 0$) is typical for generic initial states. In this regime, the relevant relaxation timescale may be set by $1/|{\rm Re}(\lambda_{k})|$ with $k\geq 3$, instead of $1/|{\rm Re}(\lambda_{2})|$. We discuss the conditions for this phenomenon in detail in [82]. The TSME should be distinguished from the conventional SME, as it represents the intrinsic property of the underlying dynamics (Lindbladian). In contrast, the SME only reflects the property of a specific initial state.